You are selling red candy for $2 and blue candy for $1. Each person is asked to sell $10 worth of candy.
1. Write and equation that says the total amount from red candy, x and blue candy, y is EXACTLY equal to $10.... IS THIS RIGHT 2x+1y=10??
2. Rewrite the equation above in slope-intercept form? IS THIS RIGHT y=10-2x???
3. What are 6 combinations of red candy and blue candy that you must sell to meet or exceed the $10 requirement..... IS THIS RIGHT 5,1 5,2 5,3 5,4 5,5 5,6????

Answer by ankor@dixie-net.com(22740)

You are selling red candy for $2 and blue candy for $1. Each person is asked to sell $10 worth of candy.
1. Write and equation that says the total amount from red candy, x and blue candy, y is EXACTLY equal to $10.... IS THIS RIGHT 2x+1y=10??
Yes this is right
:
2. Rewrite the equation above in slope-intercept form? IS THIS RIGHT y=10-2x???
Yes but usually written: y = -2x + 10 (the y = mx+b) form)
:
3. What are 6 combinations of red candy and blue candy that you must sell to meet or exceed the $10 requirement..... IS THIS RIGHT 5,1 5,2 5,3 5,4 5,5 5,6????
If you just want to meet the $10 requirement:
y = -2x + 10
If you have 5 red candies, x=5, you would have 0 blue candies (y)
y = -2(5) + 10
y = -10 + 10
y = 0
:
A table for this would look like this starting with x=0, (no red candy)
x | y
-------
0 |10
1 | 8 (1 red, 8 blue)
2 | 6
3 | 4
4 | 2
5 | 0 (5 red no blue)
Note that 6 combinations is all you, could have (no negative candy)